# In-Context Learning Strategies Emerge Rationally

### v1

@@ Abstract
Recent work analyzing in-context learning (ICL) has identified a broad set of strategies that describe model behavior in different experimental conditions. We aim to unify these findings by asking why a model learns these disparate strategies in the first place. Specifically, we start with the observation that when trained to learn a mixture of tasks, as is popular in the literature, the strategies learned by a model for performing ICL can be captured by a family of Bayesian predictors: a memorizing predictor, which assumes a discrete prior on the set of seen tasks, and a generalizing predictor, wherein the prior matches the underlying task distribution. Adopting the lens of rational analysis from cognitive science, where a learner’s behavior is explained as an optimal adaptation to data given computational constraints, we develop a hierarchical Bayesian framework that almost perfectly predicts Transformer next-token predictions throughout training without assuming access to its weights. Under this framework, pretraining is viewed as a process of updating the posterior probability of different strategies, and its inference-time behavior as a posterior-weighted average over these strategies’ predictions. Our framework draws on common assumptions about neural network learning dynamics, which make explicit a tradeoff between loss and complexity among candidate strategies: beyond how well it explains the data, a model’s preference towards implementing a strategy is dictated by its complexity. This helps explain well-known ICL phenomena, while offering novel predictions: e.g., we show a superlinear trend in the timescale for transition to memorization as task diversity is increased. Overall, our work advances an explanatory and predictive account of ICL grounded in tradeoffs between strategy loss and complexity.
@@ 1 Introduction
Figure 1: Why Does a Model Learn Different Strategies for Performing ICL? To answer this question, we analyze three distinct settings wherein a model is trained to learn a mixture of tasks. (a) Model Behavior Transitions Between Memorizing and Generalizing Predictors. We first make the observation that across settings, as diversity of data distribution and amount of training are increased, model behavior transitions between two Bayesian predictors: a memorizing predictor, $M$, which assumes a discrete prior over seen tasks, vs. a generalizing predictor, $G$, which assumes a continuous prior over the true distribution. This recapitulates prior results on task-diversity thresholds [25] and transient generalization [24] in a unifying language. (b) A Hierarchical Bayesian Framework Provides an Explanatory and Predictive Account of ICL. The consistency of these transitions motivates a hierarchical Bayesian model of ICL, wherein a model’s inference-time behavior is framed as a posterior-weighted interpolation of the Bayesian predictors $M$ and $G$, while pretraining is seen as a process of updating the preference (posterior probability) toward different predictors. We find that under reasonable assumptions regarding neural network learning dynamics, our framework is highly predictive of model behavior throughout training, without access to model weights (bottom panel). Additionally, our framework provides an explanatory account of ICL phenomena via a tradeoff between the loss and the complexity of learned solutions (top panel).
@@
To address the questions above, we first make the observation that in popularly studied ICL settings, where a model is trained to learn a mixture of tasks, identified strategies can be unified in the language of Bayesian inference: across three distinct settings, we show models learn solutions that behaviorally match Bayes-optimal strategies towards generalizing to the distribution of tasks seen during training vs. the true distribution from which said tasks are sampled (see Fig. 1(a)). These solutions respectively assume a discrete prior over the seen tasks (a memorizing predictor) or a continuous prior over the true distribution (a generalizing predictor), capturing previously studied setting-specific solutions (e.g., those described by Raventós et al. [25] in in-context linear regression). This observation then motivates our core contribution: we propose to understand ICL by invoking the approach of rational analysis from cognitive science [38, 39, 40, 41, 42], wherein a learner’s behavior is explained as an optimal adaptation to data, given its objective and computational constraints. In our case, building on the finding that Transformers transition between memorizing and generalizing predictors, we examine how a Bayes-optimal learner would trade off between these solutions across training and data conditions, given a simplicity bias [43, 44, 45, 46, 47] and power-law scaling of loss with dataset size [48, 49]. This yields a hierarchical Bayesian framework that casts pretraining as a process of weighing preference (posterior probability) toward different solutions based on their loss and complexity. At inference, model behavior is framed as a posterior-weighted average of these solutions (which themselves are Bayesian—hence the term “hierarchical”). Deriving a closed-form expression for next-token predictions based on this framework, we are able to almost perfectly explain model behavior throughout training without assuming access to the weights. This allows us to both explain several known ICL phenomena and draw novel predictions. Overall, we make the following contributions in this work.
@@ 2 Preliminaries: Learning a Finite Mixture of Tasks
We cast prior settings used for studying ICL [25, 20, 26] as learning a finite mixture of tasks. Specifically, settings analyzed in this work involve learning a mixture distribution $\mathcal{T}_{\text{train}}$ defined over $D$ parametrized functions $\{f(\bm{w}_{1},\cdot),\dots,f(\bm{w}_{D},\cdot)\}$, or ‘tasks’. For each function $f(\bm{w})$, the parameters $\bm{w}\in\mathbb{R}^{m}$ are sampled from a predefined distribution $\mathcal{T}_{\text{true}}$ (see Fig. 2). We call $D$ the task diversity of the mixture. Every training iteration, we randomly select a function $f(\bm{w},\cdot)\in\mathcal{T}_{\text{train}}$, and use it to generate a sequence of length $C$ (details vary by setting, see below). Batches consist of independently generated such sequences. We autoregressively train Transformers (GPT-NeoX architecture [55, 56]) for a predefined number of iterations $N$. Model performance is evaluated either on in-distribution (ID) sequences drawn from $\mathcal{T}_{\text{train}}$, or on out-of-distribution (OOD) sequences drawn from the underlying distribution $\mathcal{T}_{\text{true}}$ (see App. F for further details on architecture and method). We analyze three specific instantiations of this general formulation, detailed next.
@@ 3.1 Validating the Memorizing and Generalizing Predictors
Figure 4: Relative Distance Captures Transitions in Model Behavior. We show the relative distance between model outputs and the two predictors. Marginals report the absolute distance values (e.g., KL between model and predictor outputs for the Balls & Urns setting), holding $N$ constant (for the right plot), or $D$ constant (for the top plot), and varying the other variable (denoted with the dotted line). Across all settings, we see model behavior decomposes into two phases explained by either the memorizing or generalizing predictor. In this figure, we use context length of 128, task dimensionality of 8 and MLP width of 256 for balls and urns. Linear regression has similar parameters except context length of 32, and classification has similar parameters other than MLP width of 512.
@@
We next demonstrate the validity of the memorizing and generalizing predictors for the purpose of our analysis. Specifically, we show that as experimental conditions are varied, a model’s behavior primarily transitions between these predictors. To this end, we consider two core phenomena associated with ICL—an increase in models’ OOD performance with increasing task diversity $D$ [25, 59, 60], and the ‘forgetting’ of this ability with increasing training steps $N$—a phenomenon known as transient generalization [24, 61, 30]. We first replicate these results behaviorally in Fig. 1(a), finding that, across settings, Transformers transition between performing like the memorizing predictor vs. like the generalizing predictor (see App. H for full results). Then, we make a direct comparison by computing the distance between our trained model’s next-token predictions and the predictions of the memorizing and generalizing predictors. Specifically, let $d(.,.)$ denote a distance measure (KL-divergence or Euclidean distance), and denote the Transformer model trained from scratch via $h(.)$. Then, as a function of $D$ and $N$, we can plot a heatmap of the relative distance between the trained model and the memorizing and generalizing predictors, defined as $d_{\text{rel}}=\nicefrac{{(r+1)}}{{2}}$, where $r:=\frac{d(h,G)-d(h,M)}{d(G,M)}$. This metric evaluates to 00 vs. $1$ if the model is closer to the generalizing vs. the memorizing predictor. Results are shown in Fig. 4; see App. H.3 for absolute distances. We clearly see that increasing $D$ for fixed $N$, the model first behaves like a memorizing predictor (in red), only to eventually transition to behaving like a generalizing predictor (in blue)—illustrating task-diversity effects [25]. Meanwhile, when increasing $N$ for middle values of $D$, the model starts closer to a generalizing predictor, only to eventually give way to a memorizing predictor—illustrating transient generalization [24]. More broadly, across all tasks, we see there is a clear delineation of the model behavior into two phases of $(N,D)$, such that model behavior is best explained by either the memorizing or the generalizing predictor in a given phase. Given the optimality of these predictors on the distribution of seen tasks ($\mathcal{T}_{\text{train}}$) or the underlying distribution ($\mathcal{T}_{\text{true}}$), our analysis provides an explanation for why these predictors were observed in prior work. However, several questions remain, including, why is a generalizing strategy learned even when it leads to worse ID performance, and why does varying experimental conditions change which strategy, among the memorizing and generalizing predictor, is implemented by a model. We address these questions next.
@@ 4 Answering the Why: A Hierarchical Bayesian Account of ICL
Figure 5: Our Bayesian Model Captures Transitions Between Strategies Explaining Model Behavior. We plot the posterior probability of the memorizing predictor given by our theoretical model (Eq. 4). Across three broad experimental settings—(a) Balls & Urns, (b) Linear Regression, and (c) Classification—we find our model identifies the phases best explained by a given predictor and the boundary between them, hence capturing the transition between solutions seen in a Transformer’s training (as shown by the relative distance maps). Importantly, we find our model is highly predictive of the pretrained Transformer’s behavior (next-token predictions), across both settings used for fitting the three free parameters of our model vs. unseen ones. Max color bar value for Balls & Urns and Classification is determined by the performance of a baseline predictor that always outputs the mean of the distribution $\mathcal{T}_{\text{true}}$. In this figure, we use context length of 256, task dimensionality of 8 and MLP width of 256 for balls and urns. Linear regression and classification have similar parameters except context length of 64 and 384, respectively.
@@
  • A2: Neural networks exhibit a simplicity bias, such that the complexity of implementing a predictor $Q$ follows a prior $p(Q)=2^{-K(Q)^{\beta}}$, where $K(Q)$ denotes the estimated Kolmogorov complexity of $Q$. To accommodate the fact that we care about implementation in a Transformer, we instantiate a free parameter ($\beta$) on complexity while preserving the form of a universal prior.
A1 is merely a paraphrased version of well-known power-law scaling behaviors seen in neural network training [48, 49] and dictates how quickly observed data updates model behavior. That is, it offers a functional form for the rate at which likelihood in a posterior calculation grows, i.e., the rate of evidence accumulation. Meanwhile, A2 is a well-established inductive bias of neural networks that has has found use in interpreting neural networks in past work [44, 58, 64]. This constraint dictates a model’s preference over the two predictors (affecting the prior in a posterior calculation). The specific functional form we have chosen for modeling this constraint is inspired by the extensively studied notion of universal prior from algorithmic information theory, which defines the complexity of an algorithm as the description-length of the shortest program that implements it on a universal Turing machine [65, 66, 67]. In experiments, we follow standard practice in literature [68, 69, 70] and estimate an upper bound for Kolmogorov complexity (which is intractable to compute) by applying several compression algorithms to the code and data required for describing the predictor, and taking the minimal size in bits (see details in App. F.2).
Returning to our goal, we now consider a model trained for $N$ iterations on a task-mixture $\mathcal{T}_{\text{train}}$ of diversity $D$. The log-posterior odds of the two predictor can be defined as follows.
@@
where $\Delta K(D)^{\beta}:=\log 2\cdot(K_{M}(D)^{\beta}-K_{G}(D)^{\beta})$ is the difference between the exponentiated Kolmogorov complexity of the two predictors; $\Delta L(D):=L_{G}(S_{\mathcal{T}_{\text{train}}}(\infty,D))-L_{M}(S_{\mathcal{T}_{\text{train}}}(\infty,D))$ is the difference between the average loss of the two predictors on a dataset of sequences sampled from $\mathcal{T}_{\text{train}}$; and $\gamma$ is a constant related to the term $A$ from constraint A1. To get the posterior probabilities for $M$ and $G$, we simply convert $\eta$ via the sigmoid function, denoted $\sigma(\cdot)$, yielding:
@@
We now check whether our model accurately captures the behavior of the pretrained Transformer and reproduces ICL’s phenomenology. As shown in Fig. 5, our Bayesian model yields an almost perfect prediction of Transformer next-token predictions for both seen / unseen settings. Moreover, without fitting to the relative distance maps, we find an almost perfect match between the posterior probabilities of the memorizing predictor given by our model and relative distance values. These results are replicated across 72 different maps in App. H.4 with varying MLP width, context length and task dimensionality, yielding robust support for our model. Across all maps, we find our model is highly predictive of Transformer next-token predictions, with mean $R^{2}$ of $0.97$ in Linear Regression, mean agreement of $0.92$ in Classification, and mean Spearman rank correlation of $0.97$ in Balls & Urns. Additionally, we find very strong correlations of $0.99,0.98,0.99$ between our model’s posterior probabilities and the relative distance values given by the Transformer in the Linear Regression, Classification, and Balls & Urns settings, respectively. Finally, we also examine ablations of our functional form in App. I, showing the computational constraints we assume are necessary to the success of our framework.

### v2

@@ Abstract
Recent work analyzing in-context learning (ICL) has identified a broad set of strategies that describe model behavior in different experimental conditions. We aim to unify these findings by asking why a model learns these disparate strategies in the first place. Specifically, we start with the observation that when trained to learn a mixture of tasks, as is popular in the literature, the strategies learned by a model for performing ICL can be captured by a family of Bayesian predictors: a memorizing predictor, which assumes a discrete prior on the set of seen tasks, and a generalizing predictor, where the prior matches the underlying task distribution. Adopting the normative lens of rational analysis, where a learner’s behavior is explained as an optimal adaptation to data given computational constraints, we develop a hierarchical Bayesian framework that almost perfectly predicts Transformer next-token predictions throughout training—without assuming access to its weights. Under this framework, pretraining is viewed as a process of updating the posterior probability of different strategies, and inference-time behavior as a posterior-weighted average over these strategies’ predictions. Our framework draws on common assumptions about neural network learning dynamics, which make explicit a tradeoff between loss and complexity among candidate strategies: beyond how well it explains the data, a model’s preference towards implementing a strategy is dictated by its complexity. This helps explain well-known ICL phenomena, while offering novel predictions: e.g., we show a superlinear trend in the timescale for transitioning from generalization to memorization as task diversity increases. Overall, our work advances an explanatory and predictive account of ICL grounded in tradeoffs between strategy loss and complexity.
@@ 1 Introduction
Figure 1: Why Does a Model Learn Different Strategies for Performing ICL? To answer this question, we analyze three distinct settings where a model is trained to learn a mixture of tasks. (a) Model Behavior Transitions Between Memorizing and Generalizing Predictors. We first make the observation that across settings, as diversity of data distribution and amount of training are increased, model behavior transitions between two Bayesian predictors: a memorizing predictor, $M$, which assumes a discrete prior over seen tasks, and a generalizing predictor, $G$, which assumes a continuous prior over the true distribution. This recapitulates prior results on task-diversity thresholds [25] and transient generalization [24] in a unifying language. (b) A Hierarchical Bayesian Framework Provides an Explanatory and Predictive Account of ICL. The consistency of these transitions motivates a hierarchical Bayesian model of ICL, where a model’s inference-time behavior is framed as a posterior-weighted interpolation in the Bayesian predictors $M$ and $G$, while pretraining is seen as a process of updating the preference (posterior probability) toward different predictors. We find that under reasonable assumptions regarding neural network learning dynamics, our framework is highly predictive of model behavior throughout training, without access to model weights (bottom panel). Additionally, our framework provides an explanatory account of ICL phenomena via a tradeoff between the loss and the complexity of learned solutions (top panel).
@@
To address the questions above, we first make the observation that in popularly studied ICL settings, where a model is trained to learn a mixture of tasks, identified strategies can be unified in the language of Bayesian inference: across three distinct settings, we show models learn solutions that behaviorally match Bayes-optimal strategies towards generalizing to the distribution of tasks seen during training vs. the true distribution from which said tasks are sampled (see Fig. 1(a)). These solutions respectively assume a discrete prior over the seen tasks (a memorizing predictor) or a continuous prior over the true distribution (a generalizing predictor), capturing previously studied setting-specific solutions (e.g., those described by Raventós et al. [25] in in-context linear regression). This observation then motivates our core contribution: we propose to understand ICL by invoking the approach of rational analysis from cognitive science [38, 39, 40, 41, 42], where a learner’s behavior is explained as an optimal adaptation to data, given its objective and computational constraints. In our case, building on the finding that Transformers transition between memorizing and generalizing predictors, we examine how a Bayes-optimal learner would trade off between these solutions across training and data conditions, given a simplicity bias [43, 44, 45, 46, 47, 48] and power-law scaling of loss with dataset size [49, 50]. This yields a hierarchical Bayesian framework that casts pretraining as a process of weighing preference (posterior probability) toward different solutions based on their loss and complexity. At inference, model behavior is framed as a posterior-weighted average of these solutions (which themselves are Bayesian—hence the term “hierarchical”). Deriving a closed-form expression for next-token predictions based on this framework, we are able to almost perfectly explain model behavior throughout training without assuming access to the weights. This allows us to both explain several known ICL phenomena and draw novel predictions. Overall, we make the following contributions in this work.
@@ 2 Preliminaries: Learning a Finite Mixture of Tasks
We cast prior settings used for studying ICL [25, 20, 26] as learning a finite mixture of tasks. Specifically, settings analyzed in this work involve learning a mixture distribution $\mathcal{T}_{\text{train}}$ defined over $D$ parametrized functions $\{f(\bm{w}_{1},\cdot),\dots,f(\bm{w}_{D},\cdot)\}$, or ‘tasks’. For each function $f(\bm{w})$, the parameters $\bm{w}\in\mathbb{R}^{m}$ are sampled from a predefined distribution $\mathcal{T}_{\text{true}}$ (see Fig. 2). We call $D$ the task diversity of the mixture. Every training iteration, we randomly select a function $f(\bm{w},\cdot)\in\mathcal{T}_{\text{train}}$, and use it to generate a sequence of length $C$ (details vary by setting, see below). Batches consist of independently generated sequences. We autoregressively train Transformers (GPT-NeoX architecture [56, 57]) for a predefined number of iterations $N$. Model performance is evaluated either on in-distribution (ID) sequences drawn from $\mathcal{T}_{\text{train}}$, or on out-of-distribution (OOD) sequences drawn from the underlying distribution $\mathcal{T}_{\text{true}}$ (see App. F for further details on architecture and method). We analyze three specific instantiations of this general formulation, detailed next.
@@ 3.1 Validating the Memorizing and Generalizing Predictors
Figure 4: Relative Distance Captures Transitions in Model Behavior. We show the relative distance between model outputs and the two predictors. Marginals report the absolute distance values (e.g., symmetrized KL between model and predictor outputs for the Balls & Urns setting), holding $N$ constant (for the right plot), or $D$ constant (for the top plot), and varying the other variable (denoted with the dotted line). Across all settings, we see model behavior decomposes into two phases, explained by either the memorizing or generalizing predictor. In this figure, we use context length of 128, task dimensionality of 8, and MLP width of 256 for balls and urns. Linear regression has similar parameters except context length of 32, and classification has similar parameters other than MLP width of 512.
@@
We next demonstrate the validity of the memorizing and generalizing predictors for the purpose of our analysis. Specifically, we show that as experimental conditions are varied, a model’s behavior primarily transitions between these predictors. To this end, we consider two core phenomena associated with ICL—an increase in models’ OOD performance with increasing task diversity $D$ [25, 60, 61], and the ‘forgetting’ of this ability with increasing training steps $N$—a phenomenon known as transient generalization [24, 62, 30]. We first replicate these results behaviorally in Fig. 1(a), finding that, across settings, Transformers transition between performing like the memorizing predictor vs. like the generalizing predictor (see App. H for full results). Then, we make a direct comparison by computing the distance between our trained model’s next-token predictions and the predictions of the memorizing and generalizing predictors. Specifically, let $d(.,.)$ denote a distance measure (symmetrized KL-divergence or Euclidean distance), and denote the Transformer model trained from scratch via $h(.)$. Then, as a function of $D$ and $N$, we can plot a heatmap of the relative distance between the trained model and the memorizing and generalizing predictors, defined as $d_{\text{rel}}=\nicefrac{{(r+1)}}{{2}}$, where $r:=\frac{d(h,G)-d(h,M)}{d(G,M)}$. This metric evaluates to 00 vs. $1$ if the model is closer to the generalizing vs. the memorizing predictor. Results are shown in Fig. 4; see App. H.3 for absolute distances. We clearly see that increasing $D$ for fixed $N$, the model first behaves like a memorizing predictor (in red), only to eventually transition to behaving like a generalizing predictor (in blue)—illustrating task-diversity effects [25]. Meanwhile, when increasing $N$ for middle values of $D$, the model starts closer to a generalizing predictor, only to eventually give way to a memorizing predictor—illustrating transient generalization [24]. More broadly, across all tasks, we see there is a clear delineation of the model behavior into two phases of $(N,D)$, such that model behavior is best explained by either the memorizing or the generalizing predictor in a given phase. Given the optimality of these predictors on the distribution of seen tasks ($\mathcal{T}_{\text{train}}$) or the underlying distribution ($\mathcal{T}_{\text{true}}$), our analysis provides an explanation for why these predictors were observed in prior work. However, several questions remain, including why a generalizing strategy is learned even when it leads to worse ID performance, and why does varying experimental conditions change which strategy, among the memorizing and generalizing predictors, is implemented by a model. We address these questions next.
@@ 4 Answering the Why: A Hierarchical Bayesian Account of ICL
Figure 5: Our Bayesian Model Captures Transitions Between Strategies Explaining Model Behavior. We plot the posterior probability of the memorizing predictor given by our theoretical model (Eq. 4). Across three broad experimental settings—(a) Balls & Urns, (b) Linear Regression, and (c) Classification—we find our model identifies the phases best explained by a given predictor and the boundary between them, hence capturing the transition between solutions seen in a Transformer’s training (as shown by the relative distance maps). Importantly, we find our model is highly predictive of the pretrained Transformer’s behavior (next-token predictions) across conditions used for fitting the three free parameters of our model and unseen ones. Max color bar value for Balls & Urns and Classification is determined by the performance of a baseline predictor that always outputs the mean of the distribution $\mathcal{T}_{\text{true}}$. In this figure, we use context length of 256, task dimensionality of 8 and MLP width of 256 for balls and urns. Linear regression and classification have similar parameters except context length of 64 and 384, respectively.
@@
  • A2: Neural networks exhibit a bias toward simpler solutions. Specifically, using $K(Q)$ to denote the Kolmogorov complexity for predictor $Q$, we accommodate the Transformer-specific implementation cost by defining $\tilde{K}_{\text{T}}(Q)=K(Q)^{\beta}$. Then, taking the form of a universal prior, the prior probability of learning a predictor $Q$ is $p(Q)\;\propto\;2^{-\tilde{K}_{\text{T}}(Q)}=2^{-K(Q)^{\beta}}$.
A1 is merely a paraphrased version of well-known power-law scaling behaviors seen in neural network training [49, 50] and dictates how quickly observed data updates model behavior. That is, it offers a functional form for the rate at which likelihood in a posterior calculation grows, i.e., the rate of evidence accumulation. Meanwhile, A2 is a well-known inductive bias of neural networks  [43, 44, 45, 46, 47, 48]. Our specific functional form for the prior is grounded in algorithmic information theory: Kolmogorov complexity $K(Q)$ is the length of the shortest program on a universal Turing machine that implements $Q$, and the coding theorem relates it to probability via $p(Q)\propto 2^{-K(Q)}$ [65, 66, 67]. Following common practice [68, 69, 70, 71, 44, 45], we estimate Kolmogorov complexity (which is uncomputable) as $\hat{K}$ via lossless compression: we apply several lossless compressors to the code and data for $Q$, and take the smallest resulting size (App. F.2). For brevity, we write $\hat{K}$ as $K$ below.
Returning to our goal, we now consider a model trained for $N$ iterations on a task-mixture $\mathcal{T}_{\text{train}}$ of diversity $D$. The log-posterior odds of the two predictors can be defined as follows.
@@
where $\Delta K(D)^{\beta}:=\log 2\cdot(K_{M}(D)^{\beta}-K_{G}(D)^{\beta})$ is the difference between the exponentiated Kolmogorov complexity of the two predictors; $\Delta L(D):=\overline{L_{G}}(S_{\mathcal{T}_{\text{train}}}(D))-\overline{L_{M}}(S_{\mathcal{T}_{\text{train}}}(D))$ is the difference between the average loss of the two predictors on a dataset of sequences sampled from $\mathcal{T}_{\text{train}}$; and $\gamma$ is a constant related to the term $A$ from constraint A1. To get the posterior probabilities for $M$ and $G$, we simply convert $\eta$ via the sigmoid function, denoted $\sigma(\cdot)$, yielding:
@@
We now check whether our model accurately captures the behavior of the pretrained Transformer and reproduces ICL’s phenomenology. As shown in Fig. 5, our Bayesian model yields an almost perfect prediction of Transformer next-token predictions for both seen / unseen settings. Moreover, without fitting to the relative distance maps, we find an almost perfect match between the posterior probabilities of the memorizing predictor given by our model and relative distance values. These results are replicated across 72 different maps in App. H.4 with varying MLP width, context length, and task dimensionality, yielding robust support for our model. Across all maps, we find our model is highly predictive of Transformer next-token predictions, with a mean $R^{2}$ of $0.97$ in Linear Regression, a mean agreement of $0.92$ in Classification, and a mean Spearman rank correlation of $0.97$ in Balls & Urns. Additionally, we find very strong correlations of $0.99,0.98,0.99$ between our model’s posterior probabilities and the relative distance values given by the Transformer in the Linear Regression, Classification, and Balls & Urns settings, respectively. Finally, we also examine ablations of our functional form in App. I, showing the computational constraints we assume are necessary for the success of our framework.